The set of all the points (x, y,z) satisfying the equation x−z = z−y represents a
line. true or false ?
1
Expert's answer
2020-02-25T09:55:36-0500

The equation "x-z=z-y" can be rewritten as "x + y - 2z = 0", but it is an equation of the plane instead of the line. Hence, the initial statement of the question is false.

Answer: false.